Warricks watch gains one minute every 4 hurs. He sets his Watch to the correct time at 8 am on Monday as he is going to work.
What is the real time when his watch shows 8pm on Wednesday?

Answer by lwsshak3(11628)

Warricks watch gains one minute every 4 hurs. He sets his Watch to the correct time at 8 am on Monday as he is going to work.
What is the real time when his watch shows 8pm on Wednesday?
**
8 am, Monday to 8 pm Wednesday=24+24+12=60 hrs
minutes watch gained=60/4=15 min
real time=8 pm-15 min=7:45 pm, Wednesday